A river flows at 2m/s. What is the speed throught the water of a boat that can go twice as fast doenstream as upstream?
please let me know the answer as I have no clue as to what it means... thanks!

Answer by edjones(8007)

s=speed of boat in calm water.
s+2=2(s-2)
s+2=2s-4
-s=-6
s=6 m/s
.
Ed